State, with reasons, whether the following statements are True or False.
i) Diode is a bilateral device.
ii) Parallel LCR circuit can be used as a band-pass filter.
iii) A pentavalent element is added to silicon to obtain an p-type semiconductor.
iv) Feedback in a circuit always reduces its gain.
v) Barkhausen criterion gives the condition for stabilization of an amplifier gain.
vi) When input is given to an op-amp in open loop condition the output voltage is
infinite.
vii) Rectifier in a power supply regulates its output voltage.
viii) According to De Morgens’ theorem
A(BAR) x B(BAR) = AxB(BAR)
ix) Ring counter with 4 flip flops can count 32 pulses.
x) The sensitivity of flux meter is independent of the number of turns of the search
coil.
1
Expert's answer
2020-04-26T18:51:32-0400

i) Diode is a bilateral device. False because it's a nonlinear device.

ii) Parallel LCR circuit can be used as a band-pass filter. True because a parallel RLC circuit cuts a certain frequency.

iii) A pentavalent element is added to silicon to obtain an p-type semiconductor. False because Silicon has valency of 4 and adding a pentavalent element (with 5 electrons) will lead to excess of electrons and make it an n-type semiconductor.

iv) Feedback in a circuit always reduces its gain. True because it takes a part of signal away from the output.

v) Barkhausen criterion gives the condition for stabilization of an amplifier gain. False because it gives the condition for amplitude and phase balances.

vi) When input is given to an op-amp in open loop condition the output voltage is infinite. True because it has infinite gain in open loop.

vii) Rectifier in a power supply regulates its output voltage. False. It only rectifies the signal.

viii) According to De Morgens’ theorem"\bar{A}\cdot\bar{B}=\overline{A\cdot B}". False since "\bar{A}+\bar{B}=\overline{A\cdot B}."

ix) Ring counter with 4 flip flops can count 32 pulses. False because such counter can count 8 pulses.

x) The sensitivity of flux meter is independent of the number of turns of the search coil. False because the more turns we have, the more the signal can be amplified.